A metal alloy that is 17% copper to make jewelry. How many ounces of a 14% alloy must be mixed with a 23% alloy to form 72 ounces of the desired alloy?
Answer by checkley71(8403)

.17*72=.23x+.14(72-x)
12.24=.23x+10.08-.14x
12.24=.09x+10.08
.09x=12.24-10.08
.09x=2.16
x=2.16/.09
x=24 ounces of 23% alloy is needed.
72-24=48 ounces of 14% is also needed.
proof
12.24=.23*24+.14*48
12.24=5.52+6.72
12.24=12.24